OpenStudy (awkwardpanda):

4(x+3)≤44 4x+12≤44 do i add 12 to both sides?

OpenStudy (awkwardpanda):

So i subtract because it's addition in the original problem, right?

OpenStudy (charlotte123):

@awkwardpanda yes ^-^

OpenStudy (awkwardpanda):

YAY. PROGRESSION

OpenStudy (charlotte123):

wait are these 2 equations or just one? XD

OpenStudy (charlotte123):

bahahahah XD

OpenStudy (awkwardpanda):

juST ONE XD the second one was my step from the first one. i simplified it. xD so now does the 4x become negative?

OpenStudy (charlotte123):

@awkwardpanda my bad XD lets see :3

OpenStudy (awkwardpanda):

4x+12-12 less than or equal to 44-12

OpenStudy (texaschic101):

4(x + 3) <= 44 4x + 12 <= 44 (subtract 12 from both sides) 4x <= 44 - 12 4x <= 32 x <= 8

OpenStudy (charlotte123):

@texaschic101 got it! XD

OpenStudy (awkwardpanda):

OH! HAY THANKS @texaschic101

OpenStudy (texaschic101):

no problem :)

OpenStudy (charlotte123):

The graph would look like that ^-^

OpenStudy (texaschic101):

yes it would :)

OpenStudy (charlotte123):

@texaschic101 XD thanks for the help on her question ^-^ helped me out too, now I could eat lunch :D

OpenStudy (texaschic101):

lol.......anytime. Have a good lunch

OpenStudy (charlotte123):

@texaschic101 thank u ^-^ take care XD

OpenStudy (charlotte123):

@awkwardpanda take care ヽ(*・ω・)ノ

OpenStudy (awkwardpanda):

Wait. so what makes problems 14-4y greater than or equal to 38 (answer is y less than or equal to -6) different from this one? why is it positive? sorry for the bother D:

OpenStudy (charlotte123):

@awkwardpanda your not doing anything with a negative, like multiplying or divinding it ^-^

OpenStudy (texaschic101):

14 - 4y >= 38 -->subtract 14 from both sides) -4y >= 38 - 14 -4y >= 24 (divide by -4) y <= -6 (you have to change the inequality sign when multiplying/dividing by a negative)

OpenStudy (charlotte123):

@texaschic101 exactly ^-^

OpenStudy (charlotte123):

thank u XD

OpenStudy (charlotte123):

now I shall go eat lunch ~ ^-^ take care everyone, and thanks again @texaschic101 ^O^

OpenStudy (texaschic101):

no problem :)